OpenStudy (anonymous):

Evaluate the given expressions. All numbers are approximate. 15.66^2-(-4.017)^4/1.044(-3.68) how do i do this?

OpenStudy (anonymous):

@campbell_st i have to do it by significant digits to gethe correct answers

OpenStudy (campbell_st):

well type it into a calculator as \[15.66^2 - \frac{(-4.017)^4}{(1.044 \times -3.68)}\]

OpenStudy (wolf1728):

campbell_st that looks a LOT better that way

OpenStudy (campbell_st):

so once you get an answer, then identify the significant digits

OpenStudy (anonymous):

i did that and i got -15.14422269/-3.84192 and this equals 3.941

OpenStudy (anonymous):

but not sure if its right

OpenStudy (campbell_st):

ummm I got 313.0089...

OpenStudy (anonymous):

cause my teacher had one similar to this 1328^2-(-7.023)^4/1.037(-2.01) = -2256/-2.08437= 1082.341427=1080 i guess do to signifigant digits

OpenStudy (wolf1728):

Yes Ijust got the same answer 313.0089588132

OpenStudy (wolf1728):

Ash90 - Is the way you typed that exactly the same way as campbell_st retyped it?

OpenStudy (anonymous):

let me double check

OpenStudy (campbell_st):

so lets assume 313.0089... is the answer... how many significant figures do you need to answer to?

OpenStudy (anonymous):

i didn't get that. but if 131.0080 is the answer you would need 4 signifigants correct?

OpenStudy (wolf1728):

Ash90 - provided that campbell_st didn't misunderstand what you had typed

OpenStudy (anonymous):

sorry i get 131.6048806

OpenStudy (wolf1728):

Or me for that matter

OpenStudy (anonymous):

but you would need 4 sigs right?

OpenStudy (campbell_st):

well for significant figures... the rule is find the 1st non zero digit... then count form there in your example its 1 then count 4 digits its 131.6 look at the digit after the 6 and apply the rounding rule... 0 so no change so the answer would be 131.6

OpenStudy (wolf1728):

but you need the answer correct too

OpenStudy (anonymous):

yes i do wolf and campbell i yes it would be131.6 i see how you got that

OpenStudy (wolf1728):

So, Ash90 is this how you are doing your calculations?

OpenStudy (anonymous):

yes

OpenStudy (wolf1728):

Well that would expxplain a LOT

OpenStudy (anonymous):

yeah

OpenStudy (wolf1728):

Want to work that out right here? I'll do the posting.

OpenStudy (anonymous):

correct

OpenStudy (wolf1728):

15.66²=245.2356

OpenStudy (wolf1728):

-4.017^4 = 260.3798226915

OpenStudy (anonymous):

4.017^4= yes correct

OpenStudy (anonymous):

so far so good

OpenStudy (wolf1728):

1.044 * (-3.68) = -3.84192

OpenStudy (anonymous):

yes i have that too

OpenStudy (wolf1728):

Okay so we have (245.2356 - 260.3798226915) / -3.84192

OpenStudy (anonymous):

245.2356-260.3798226915=15.14422269

OpenStudy (wolf1728):

I get + 15.1442269

OpenStudy (anonymous):

-15.14422269/-3.84192=3.941837074

OpenStudy (anonymous):

the answer will be 3.941837074 or a negative of that number right?

OpenStudy (wolf1728):

I get 3.941837074 So, we agree on the final outcome.By the way I notice you automatically change 15.1442269 to negative 15.1442269 in the next step. By the way it is definitely a positive number3.941837074

OpenStudy (wolf1728):

Gee I wish these posts were numbered - it would be so much easier to say - 'see post 27'

OpenStudy (anonymous):

haha lol and thanks a lot

OpenStudy (anonymous):

ans so for the sig we get 3.942 correct

OpenStudy (wolf1728):

By the way, here's the post where you have the positive number 15.14422269 245.2356-260.3798226915=15.14422269 (You turn it negative later on). 3.941837074 = 3.942 Yes, that would be it to 4 significant figures

OpenStudy (anonymous):

okay gotcha later on switch the sign

OpenStudy (wolf1728):

Okay but the answer is definitely positive.

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

thanks

OpenStudy (wolf1728):

u r welcome